OpenStudy (anonymous):

im trying to prove the inverse triangle inequality\[||x|-|y||\leq|x-y|\]like this:\[x=x-y+y,\]\[|x|=|x-y+y|,\]\[|x|\leq|x-y|+|y|,\]\[|x|-|y|\leq|x-y|.\]i get stuck here; dont know if its legit or not to take the absolute value of the LHS and preserve the inequality. help?

OpenStudy (watchmath):

your statement is equivalent to \(-|x-y|\leq |x|-|y|\leq |x-y|\) now use triangle inequality

OpenStudy (anonymous):

im a little confused... how can i do that?

OpenStudy (watchmath):

Do you know that the inequality \(|x|\leq a\) is equivalent to \(-a\leq x\leq a\) ?

OpenStudy (anonymous):

i agree, but how can i apply the triangle inequality to the above system?

OpenStudy (anonymous):

i thought i could only apply it to an expression.

OpenStudy (watchmath):

ok, let's look at the first inequality \(-|x-y|\leq |x|-|y|\). Can you rewrite the inequality so that there is no minus (but all plus) by moving around the expressions?

OpenStudy (watchmath):

Can you see that it is equivalent to \(|y|\leq |x|+|x-y|\)?

OpenStudy (anonymous):

\[-|x-y|\leq|x|-|y|\implies|y|\leq|x|+|x-y|?\]

OpenStudy (watchmath):

great!

OpenStudy (watchmath):

Now if we apply the triangle inequality to \(|y|=|(y-x)+x|\) what do we get?

OpenStudy (anonymous):

\[|y|\leq|y-x|+|x|\]

OpenStudy (watchmath):

yes, but that is the same as \(|y|\leq |x|+|x-y|\) right ? since |y-x|=|x-y|

OpenStudy (anonymous):

i agree.

OpenStudy (watchmath):

so if you want a more linear argument we can like this for the first inequality: \(|y|=|(y-x)+x|\leq |y-x|+|x|=|x|+|x-y|\) It follows that \(-|x-y|\leq |x|-|y|\)

OpenStudy (watchmath):

you can do something similar to the 2nd one

OpenStudy (anonymous):

that makes sense. and with this, we can conclude that\[||x|-|y||\leq|x-y|?\]

OpenStudy (watchmath):

yes, if you prove two inequalities: \(-|x-y|\leq |x|-|y|\) and \(|x|-|y|\leq |x-y|\) you conclude the original inequality

OpenStudy (anonymous):

beautiful. i see it now. thank you very much for your time and help. i appreciate it a lot!

OpenStudy (watchmath):

you are welcome :)